OpenStudy (anonymous):

Use a clock or stopwatch to take your pulse rate while sitting relaxed. Record the number of beats for 30 seconds. Now, walk out your front door and run around your house, then come back inside. Take your pulse again and record the number of beats for 30 seconds. What is the percent increase in your pulse rate? Show your work. Round your percent to the nearest whole number.

OpenStudy (anonymous):

...its your heartrate. what was your heartrate?

OpenStudy (australopithecus):

what if you live in a condo how are you suppose to run around your house then

OpenStudy (australopithecus):

or an apartment building

OpenStudy (anonymous):

wtf idk ahahaha

OpenStudy (australopithecus):

well if you don't live in a house you can run around I would just not answer it

OpenStudy (unklerhaukus):

Have you taken some measurements yet?

OpenStudy (australopithecus):

but yeah take some measurements or something then we can help you out

OpenStudy (australopithecus):

The answer should be ((Final heart Rate)/(Initial heart Rate))*100 - 100

OpenStudy (australopithecus):

This is based on the fact that your initial heart rate will be 100%

OpenStudy (australopithecus):

You will end up with above 100%, to correct for this we subtract by 100

OpenStudy (australopithecus):

since your final heart rate will be higher than your initial, I'm explaining it to the person who asked the question :)

OpenStudy (australopithecus):

I'm not trying to rub it in lol

OpenStudy (unklerhaukus):

i took a measurement \[P_{_{resting}}=35\frac{[\text{beats}]}{30[\text s]}\times\frac{60[\text{s}]}{[\text{min}]}=70\frac{[\text{beats}]}{[\text {min}]}\]

OpenStudy (shane_b):

What if your final heart rate was lower than your initial? I mean, you could have died during the running. I wonder if this teacher has a good lawyer.

OpenStudy (australopithecus):

That wouldn't make any sense Shane_B

OpenStudy (australopithecus):

logically your heart rate at rest is always slower than your heart rate after being active

OpenStudy (shane_b):

Did you read what I said ?!

OpenStudy (australopithecus):

Combined me and rukus have answered your question :)

OpenStudy (australopithecus):

if you are still confused please feel free to ask any questions you may have @allstargirl863 @Shane_B yes I did your heart rate will always be higher after physical activity, This question assumes the individual is at rest initially.

OpenStudy (anonymous):

okay

OpenStudy (shane_b):

I think you missed the joke altogether...I'm pretty sure when you're dead, your heart rate is 0. I give up :)

OpenStudy (australopithecus):

Oh no I get it :) but no one would be around to answer the question anymore. You are right also in a case like this it would probably be the government that falls under liability not the teacher unless he has tons of money.

OpenStudy (australopithecus):

I cannot recall the legal term

OpenStudy (australopithecus):

Sorry I only half read things normally

OpenStudy (shane_b):

I have the same problem sometimes :)

OpenStudy (unklerhaukus):

i took another measurement \[P_{_\text{resting+three coffees}}=40\frac{[\text{beats}]}{30[\text s]}\times\frac{60[\text{s}]}{[\text{min}]}=80\frac{[\text{beats}]}{[\text {min}]}\]

OpenStudy (unklerhaukus):

increase \[P_{\text{resting+three coffees}}-P_{\text{resting}}\] percentage increase \[\frac{P_{\text{resting+three coffees}}-P_{\text{resting}}}{P_{\text{resting}}}\times100\%\]\[=\frac{80\frac{[\text{beats}]}{[\text {min}]}-70\frac{[\text{beats}]}{[\text {min}]}}{70\frac{[\text{beats}]}{[\text {min}]}}\times100\%\]\[=\frac 17\times100\%\]\[\approx14\%\]